A manufacturer of backpacks plans to introduce a new line. Equipment and production costs will be incurred immediately and will total $7,272,727. The company expects to earn a profit of $20 per backpack, and sales are estimated to be 100,000 in the first year (assume that the cash flow comes in at the end of the year). Sales are then expected to grow by 10% per year in each of the next three years (in year 2, year 3, and year 4) but the price is expected to remain at $20 throughout. What is the internal rate of return of this project?

______%

Place your answer in percentage form with no percentage sign. That is, if your answer is four point eight eight percent, you should enter that value as 4.88.

Should the company produce the backpacks if the required rate of return is 12%?
________(Yes or No)

You must get both parts correct to receive credit.
Answer #1

a). Profit in year 1 = Profit per backpack * Sales = $20 * 100,000 = $2,000,000

Profit in year 2 = Profit in year 1 * (1 + g) = $2,000,000 * (1 + 0.10) = $2,200,000

Profit in year 3 = Profit in year 1 * (1 + g) = $2,200,000 * (1 + 0.10) = $2,420,000

Profit in year 4 = Profit in year 1 * (1 + g) = $2,000,000 * (1 + 0.10) = $2,662,000

To find the IRR, we need to put the following values in the financial calculator:

CF0 = -7,272,727; C01 = 2,000,000; F01 = 1; C02 = 2,200,000; F02 = 1; C03 = 2,420,000; F03 = 1; C04 = 2,662,000; F04 = 1;

Press IRR, then CPT, which gives us 10

So, IRR = 10%

b). No, the company should not produce the backpacks as the required rate of return(12%) is more than the IRR(10%).